Mr A has only 4 TV channels all of them are quite boring so he natuarally desire to change channel after every one minute .The probability that he is back to origional channel for first time after 4 minutes is...

Grade:12

4 Answers

Yash Kawdiya
6 years ago
The probablity to come on 1st channel is equals to 1. as their are only four channels so we get to the first on after 4 min. So Final Ans is 1.
ANKIT
6 years ago
if A,B,C,and D are chennel , and he can change the channel in both side( assume cyclic arrenge ment , becaus he can either increase the channel no or decrease  the channel no ) so for after 4th  time  he will get the original (1st ) channel there is only one way . and now we can see he has two ways to go  so total number of ways he can change channel is 2*2*2*2=16 
P(required) 1/16. 
sarayu
2 years ago
The probability that he is back to origional channel for first time after 4 minutes is
p=3/2*2/3*2/3*1/3=1/3
 
sarayu
2 years ago
The probability that he is back to origional channel for first time after 4 minutes is...
p=3/2*2/3*2/3*1/3=1/3